OpenStudy (anonymous):

An automobile maker has an order for 7,500 new cars to be delivered in one week. Each car must be fitted with a new hood ornament that requires 1 hour and 15 minutes to install. Assuming the factory operates on a standard workweek shift (8 hours per day for 5 days), how many workers must be assigned this job to meet the deadline?

OpenStudy (anonymous):

I got it. It is 235 workers.

OpenStudy (anonymous):

time to fit 7500 cars = 7500 * 1.25 hours number of hours available in week = 40 hours number of workers needed = 7500* 1.25 / 40 = 234.4 number required = 235

OpenStudy (anonymous):

But thank you

OpenStudy (anonymous):

alright - great minds think alike - lol

OpenStudy (anonymous):

:) haha